OpenStudy (anonymous):

Evaluate the function f(x) = 3x − 2, when x = −1. −9 −5 1 5

OpenStudy (anonymous):

f(x) = 3x − 2, when x = −1. Replace x by -1 f(-1)=-1*3-2=-3-2=-5